OpenStudy (anonymous):

I need help to figure out how to solve this problem. Determine whether each pair of equations represents parallel lines. y=5-3x 3x-y=8

jimthompson5910 (jim_thompson5910):

what do you get when you solve 3x-y=8 for y

OpenStudy (anonymous):

You know that lines are parallel if they have equal slopes so put both equations in y=mx+b form in order to compare their slopes (m).

OpenStudy (anonymous):

would it be y=3?

OpenStudy (anonymous):

What do you mean by y=3?

OpenStudy (anonymous):

I used the formula y=mx+b like Twis7ed said. I may have done it wrong but that is the answer i came up with.

jimthompson5910 (jim_thompson5910):

solve 3x-y=8 for y to get 3x-y=8 3x-y-3x=8-3x -y = 8 - 3x -y = -3x + 8 y = 3x - 8

jimthompson5910 (jim_thompson5910):

Now compare slopes

jimthompson5910 (jim_thompson5910):

Keep in mind that y=5-3x is the same as y=-3x+5

OpenStudy (anonymous):

I am even more confused now. Is the 3x the point on the graph on the x axis and the 5 the point on the y axis?

jimthompson5910 (jim_thompson5910):

y=5-3x turned into y=-3x+5 The slope of this is -3 (since that's the number in front of the x) ------------------------------------------------------- 3x-y=8 turned into y = 3x - 8 which has a slope of 3

jimthompson5910 (jim_thompson5910):

So what can you tell me about these two lines if you know the slopes are -3 and 3

OpenStudy (anonymous):

The lines are not parallel

jimthompson5910 (jim_thompson5910):

good, the two slopes would have to be equal if the lines were parallel

OpenStudy (anonymous):

Thank you. That was more helpful that the book.